# Disentangling Brillouin’s negentropy law of information and Landauer’s law on data erasure

### v2

@@ II.3 Information theory: come back of the observer
    In no case the average number of bits per character can be less than
@@
    $H$ is named quantity of information emitted by the source and by identification with Eq.1, $S=H\times\ln 2$ is its entropy.
@@
    Within a factor, $H$ (and thus $S$) is the only measure of uncertainty on the upcoming character that is 1) continuous in $p$; 2) increasing in $\mathcal{W}=1/p$ for uniform distributions; 3) additive over different independent sources of uncertainty.
@@
These results are general and apply regardless of the random variable, and in particular to the constantly evolving phases of any physical dynamic system. It follows that, having been a quantity of heat dissipated (Clausius), then the logarithm of a number of possibilities (Gibbs), entropy is now (Shannon), and all at once, the uncertainty that the observer has concerning the phase of the system. Here again, its understanding had made great progress.
@@
These two results were quickly recognized as a major advance for thermostatics. Brillouin, with his law of information [6, 7] and Jaynes, with the maximum entropy principle [33, 34], were the firsts. The first point will be presented in the following section; we will therefore give a brief overview of the second here.
@@
How to describe or mentally represent, in a rational way, something that we only partially know? That is in reality the central question of science. The description must account for all pieces of information, otherwise it would be incomplete, but it must not invent an information that comes out of nowhere (which would constitute a synthetic a priori knowledge). For instance, this is the problem of linear-regression: for the solution (the description) to be unique, the degree of the polynomial must be less than the number of points. Solving such a problem amounts to seeking a unique description that maximizes the uncertainty about the information we do not have. For a description involving a probability distribution, Shore and Johnson [35] showed that maximizing its entropy (instead of another possible measure of uncertainty) is the only procedure leading to a unique solution.
@@
Hence the theorem of maximum entropy: the best prior probability distribution that accounts for our knowledge is that of maximum entropy. For instance, Shannon [4] showed that “when one does not know anything…” (Balian) the maximum of entropy is obtained for a uniform distribution. This provides a mathematical basis for Laplace’s principle of insufficient reason and makes it analytical.
@@
By getting ride of the ergodic hypothesis, there is absolutely no longer any obstacle, no inconsistency, in postulating by convention that equilibrium is an attractor, that is to say, in defining equilibrium as the state of maximum entropy of phases distribution (that is also the state of maximum entropy of variables whose distributions are similarity-invariant in form [36], such as the density). This is the principle of maximum entropy [33, 34]. This principle, like any other definition, is conventional. It is not required to be checked by experiment. And thankfully so, because it would not be possible. This principle is just convenient. Without introducing inconsistency, it allows us to establish a deductive link between theoretical statements (the definition of equilibrium plus the 2nd law) and experiments (the observed restoring force towards equilibrium), that is to say, to explain and account for the latter.
@@ III.1 From Maxwell to Szilard
The connection between information and energy was established by Maxwell [3]. He imagines an intelligent being (say a demon or a computer) capable of tracking particles and, from this knowledge, extracting energy from the system which would otherwise be impossible. A simpler version of such an engine is that of Szilard [37], shown in Figure 2. Applying the first law of thermodynamics, which states that energy cannot by definition come from nowhere, forces us either to admit that information is a form of energy or to formulate another definition of energy. Undeniably, the first alternative is more convenient.
@@
Figure 2: Cycle of operations of Szilard’s engine. By knowing the location of the particle, a demon (say a computer) can produce mechanical work with no other cost than that of acquiring the information. The first principle implies that information is energy.
@@ III.2 Brillouin’s negentropy law of information
    Clausius: The negative of the entropy difference $\Delta S$ experienced by a system (at given $T$) is the minimum work $W$ that must be done on the system for the change to take place: $W\geq-T\Delta S$.
@@
    Shannon: The Gibbs formula for $S$ is actually to a factor $\ln 2$ that of the uncertainty $H$ on the actual microstate: $S=H\ln 2$
@@
    Brillouin: Therefore, reducing uncertainty by acquiring information requires minimum work: $W\geq-T\Delta H\ln 2$. For one bit of acquired information $\Delta H=-1$, so that:
@@
where $W_{\text{acq/bit}}$ is the minimum work that we have to provide (and that will be dissipated as heat) per bit of acquired information. This is the Brillouin’s negentropy law of information [6, 7], called “principle” by Brillouin [6], also called sometime Szilard’s principle [38] (because Eq.4 can be derived from the functioning of Szilard engine), but here called “law” for the sake of consistency with §II.1 (because it is derived from the 2nd law). With Brillouin’s equation, the energy balance of Szilard engine is zero, as required.
Let us consider a body of mass $m$, lifted by a height $\Delta h$. Classical mechanics tell us that the increase in potential energy is $mg\Delta h$. This result does not impose any particular steps for the change in height and does not specify where the mechanical work is done. However, the result concerning the net change in potential energy is general and valid in all cases, regardless of the path taken.
Brillouin’s law works in exactly the same way. Entropy, and consequently the uncertainty about a system or the information we lack to describe it, is a state quantity whose variation is independent of the path connecting two states. Thus, Brillouin’s law does not tell us where the energy cost of acquisition is paid in the operation. And this is inherent to the definition of what is a state quantity. Acquisition may not be a simple operation, but may consist of other more elementary operations, which may differ depending on the specific data acquired. But at the end, whatever the way acquisition is done, it will cost at least $T\ln 2$ in average per bit.
Brillouin’s law is a syllogism and says nothing more than its premises, in this case two mathematical results and the second law. It follows that its generality is as robust as that of the second law.
@@ III.3 Landauer’s law on data erasure
Landauer’s competing argument [8] is based on two main ideas. The first is that “information is not a disembodied abstract entity; it is always tied to a physical representation. It is represented by engraving on a stone tablet, a spin, a charge, a hole in a punched card, a mark on paper, or some other equivalent”[10]. In other words, information is always supported by a hardware with two discernible different configurations (two values 0 or 1), namely a data-bit. The second idea is that the acquisition of one bit of information necessarily passes by the erasure of its supporting data-bit. The reasoning is the following:
@@
    Any intelligent being has a finite memory, so that the infinite cyclic acquisition of information about a dynamical system necessarily requires the erasure of data-bits.
@@
    Erasing a data-bit (a thermodynamical system) consists of setting it to an arbitrary value (say 0). The procedure must be able to work for a known or unknown initial value (i.e. it must be the same in both cases). This constraint automatically implies a two-step erasure process (see Figure 3):
@@
    The first step does not involve any exchanges with the environment, whereas the second dissipates at least $T\ln 2$ of heat, or equivalently at constant internal energy (i.e. at constant temperature for a set of independent data-bits), requires a minimum work. The net balance of the two gives
@@
    which leads, as Brillouin’s law does (Eq.4), to a zero energy balance for Szilard engine. The difference lies in the fact that, according to Landauer, erasure is a necessary step and it is at this very point that the energy cost of data acquisition is paid.
Figure 3: Landauer erasure of a data-bit consisting in a single particle in a two-compartment box. The procedure must be independent of the initial position of the particle: neither a simple shift of the particle towards 0 by means of two pistons keeping constant the volume of the phase space; nor a quasi-static expansion with a single piston is possible. According to Landauer, the first step is necessarily a thermodynamically irreversible free expansion.
@@
    Restricted context of data erasure: does erasing one data-bit really require at least $T\ln 2$ (Landauer’s limit) of work?
    To my knowledge, all the authors (see e.g. [39, 40, 41, 42, 43, 44]) who have addressed this question, actually used the Landauer’s procedure that consists in: a) free expansion of the phase space; b) reversible compression of the phase space. By doing this the authors simply test the 2nd law and not the new point of Landauer’s principle, so that their results are not surprising because there is absolutely no error in the calculation made by Landauer. What is new in Landauer’s assertion? It is the claim that there is no alternative erasing procedure. Thus, the only way to confirm Landauer’s principle would be to search, in vain, for an alternative. However, this is not what was done. In reality, even within the restricted context of data erasure, several points of Landauer’s reasoning are questionable: the imperative to begin erasure with expansion, and the necessity of this expansion to be thermodynamically irreversible. Counterexamples have been provided [45, 46, 47] which invalidate its generality.
@@ IV Information versus data
The confusion between information and data is also that between the loss (or deletion) of a bit of information and the erasure of a bit of data. The first refers to the observer’s knowledge, whereas the second only involves a physical object independent of them. A similar confusion in statistical mechanics has already led to inconsistencies. It is therefore not surprising to see others with Landauer’s idea. These inconsistencies make Landauer’s law incompatible with that of Brillouin, that is to say incompatible with the triplet (thermodynamics-statistical mechanics-information theory).
@@ IV.1 Total versus incomplete information
In thermodynamics, setting a data-bit into the S state, by removing the partition between two compartments containing a single particle (1st step of Landauer erasure), is reversible or not depending on whether the initial value of the bit is known or not [48] (see figure 4). This problem is the same as that of the Gibbs paradoxes and the solution also: thermodynamics considers the case where the observer has an incomplete information about the system, whereas Landauer in his reasoning implicitly assumes that they have the full information at their disposal. There is no solution to this paradox other than those that reintroduce the observer, because in reality, taking these two cases into account (as Bennett does [49], a defender of Landauer’s thesis) already amounts to taking the observer into account.
Figure 4: Setting a data-bit in state S by merging two compartments is thermodynamically irreversible (red path) if the initial value is known ($\Delta S=\ln 2$), but it is thermodynamically reversible (green path) if the initial value is unknown ($\Delta S=0$). This is reminiscent of the Gibbs paradox of mixing. Landauer viewpoint conflicts with thermodynamics.
In addition, the constraint of a single path for erasure, regardless of the initial value of the bit, stems from the requirement to work with unknown values. But this case is thermodynamically reversible. For known values, there is nothing preventing the use of two different reversible paths. The use of a single erasing procedure for known and unknown cases is an arbitrary choice by the observer. Here again lies a clear inconsistency between the intention of theory and its result, the desire to eliminate the observer while reintroducing him.
@@ IV.2 Global concept versus local object
Information is a concept that must not be confuse with its material embodiment. The proof is given by Landauer himself: “Information […] is represented by engraving on a stone tablet, a spin, a charge, a hole in a punched card, a mark on paper, or some other equivalent”[10]. The material embodiment may change but information remains the same. The denial of this difference leads to severe inconsistencies.
@@
A fundamental property of information is that it cannot be given twice. We can duplicate bits of data encoding certain information so that the storage space occupied is twice as large as before, but the corresponding amount of information remains unchanged. Conversely, if multiple copies of the same data-bit are erased, the corresponding information will only be lost when the last copy is erased. To assert that information has been erased, one must take into account not only what happened locally, but also the existence or not of a copy somewhere in a more global space.
Information is a global concept, whereas a data-bit is a local object.
@@
It could be argued that this is playing with words and that information can have different meanings. Actually, the common meaning of information as “The imparting of knowledge in general. Knowledge communicated concerning some particular fact, subject, or event” (Oxford Dictionary); is also that of Shannon (if we know the language used for a message, we can compress it further, but indicating the language twice does not allow us to save more space); but this meaning is also that of the defenders of Landauer’s idea (“But what is information? A simple, intuitive answer is “what you don’t already know””[41]. “If someone tells you that the earth is spherical, you surely would not learn much”[43]).
@@
The denial of the difference between information and data and the attempt to materialize information is culminating with the supposed “information-mass equivalence”[50, 12, 51, 52] that leads to paradoxes in relation with the corresponding mass deficit. For instance in the framework of this equivalence, consider a data-bit encoding a bit of information. Duplicate the bit and make the original and the copy physically independent. Erase one of the two. The quantity of information remains unchanged, so that the erased data-bit does not display any mass deficit. Erase the second bit. Information is lost that corresponds to a mass deficit. How does the second bit “knows” that the first has been erased? This contradicts the hypothesis of independence. But if we suppress this hypothesis and assert that independence is impossible, does not that mean that information is not a property of the local data-bit, but a property of a global system?
@@
Yet another variation of this paradox. Consider a data-bit, encoding for a bit of information, whose initial value is 0. Erase the data-bit (set its value to 0). There is no longer information stored by the data-bit. Erase the bit again. What is the difference between the two erasures? If there is no difference, an eventual mass deficit should not be due to any loss of information. It there is a difference, how the data-bit “knows” that it has been already erased? Only a global observer could.
@@
A data-bit has a physical meaning at all scales: from that of a single particle in a two-compartment box, to the macroscopic scale. But information is meaningful only when it involves an observer, information is meaningful only at macroscale. Information, even the smallest piece, is an emergent concept.

### v3

@@ II.3 Information Theory: The Return of the Observer
    In no case, the average number of bits per character is less than
@@
    $H$ is named quantity of information emitted by the source and by identification with Equation (1), $S=H\times\ln 2$ is its entropy.
@@
    Within a factor, $H$ (and thus $S$) is the only measure of uncertainty on the upcoming character that is (1) continuous in $p$; (2) increasing in $\mathcal{W}=1/p$ for uniform distributions; and (3) additive over different independent sources of uncertainty.
@@
These results are general and apply regardless of the random variable, and in particular to the constantly evolving phases of any physical dynamic system. It follows that entropy, having first been defined as a quantity of dissipated heat (Clausius) and then as the logarithm of a number of possibilities (Gibbs), is now—according to Shannon—the uncertainty that the observer has concerning the phase of the system. Here again, its understanding had made great progress.
@@
These two results were quickly recognized as a major advance in thermostatics. Brillouin, with his law of information [5, 6] and Jaynes, with the maximum entropy principle [28, 27], were the first. The first point will be presented in the following section; therefore, we provide a brief overview of the second here.
@@
How to describe or mentally represent, in a rational way, something that we only partially know? That is, in reality, the central question of science. The description must account for all pieces of information; otherwise, it would be incomplete, but it must not invent information that comes out of nowhere (which would constitute a synthetic a priori knowledge). For instance, this is the problem of linear regression: for the solution (the description) to be unique, the degree of the polynomial must be less than the number of points. Solving such a problem amounts to seeking a unique description that maximizes the uncertainty about the information we do not have. For a description involving a probability distribution, Shore and Johnson [51] showed that maximizing its entropy (instead of another possible measure of uncertainty) is the only procedure leading to a unique solution.
@@
Hence, the theorem of maximum entropy: the best prior probability distribution that accounts for our knowledge is that of maximum entropy. For instance, Shannon [50] showed that “when one does not know anything…” (Balian) the maximum of entropy is obtained for a uniform distribution. This provides a mathematical basis for Laplace’s principle of insufficient reason and makes it analytical.
@@
By abandoning the ergodic hypothesis, there is no longer any obstacle or inconsistency in postulating by convention that equilibrium is an attractor—that is, in defining equilibrium as the state of maximum entropy of the phase distribution (which is also the state of maximum entropy of variables whose distributions are similarity-invariant in form [29], such as the density). This is the principle of maximum entropy [28, 27]. This principle, like any other definition, is conventional. It is not required to be checked by experiment. And, thankfully so, because it would not be possible. This principle is just convenient. Without introducing inconsistency, this allows us to establish a deductive link between theoretical statements (the definition of equilibrium plus the second law) and experiments (the observed restoring force towards equilibrium), that is, to explain and account for the latter.
@@ III.1 From Maxwell to Szilard
The connection between information and energy was established by Maxwell [44]. He imagines an intelligent being (say, a demon or a computer) capable of tracking particles and, from this knowledge, extracting energy from the system, which would otherwise be impossible. A simpler version of such an engine is that of Szilard [52], shown in Figure 2. Applying the first principle of thermodynamics, which states that energy cannot, by definition, come from nowhere, forces us either to admit that information is a form of energy or to formulate another definition of energy. Undeniably, the first alternative is more convenient.
Figure 2: Cycle of the operations of Szilard’s engine. By knowing the location of the particle, a demon (say a computer) can produce mechanical work with no other cost than that of acquiring the information. The first principle implies that information is energy.
@@ III.2 Brillouin’s Negentropy Law of Information
    Clausius: The negative of the entropy difference $\Delta S$ experienced by a system (at a given $T$) is the minimum work $W$ that must be performed on the system for the change to take place: $W\geq-T\Delta S$.
@@
    Shannon: The Gibbs formula for $S$ is actually to a factor of $\ln 2$ that of the uncertainty $H$ on the actual microstate: $S=H\ln 2$
@@
    Brillouin: Therefore, reducing uncertainty by acquiring information requires minimum work: $W\geq-T\Delta H\ln 2$. For one bit of acquired information $\Delta H=-1$; thus
@@
    where $W_{\text{acq/bit}}$ is the minimum work that we have to provide (and that will be dissipated as heat) per bit of acquired information. This is the Brillouin’s negentropy law of information [5, 6], called “principle” by Brillouin [5] and sometimes referred to as Szilard’s principle [13], since Equation (4) can be derived from the operation of the Szilard engine. Here, however, it is referred to as a “law” for the sake of consistency with Section II.1, because it is derived from the second law. With Brillouin’s equation, the energy balance of the Szilard engine is zero, as required.
Let us consider a body of mass $m$, lifted to a height $\Delta h$. Classical mechanics tells us that the increase in potential energy is $mg\Delta h$. This result does not impose any particular steps for the change in height and does not specify where the mechanical work is performed. However, the result concerning the net change in potential energy is general and valid in all cases, regardless of the path taken.
Brillouin’s law works in exactly the same way. Entropy, and consequently the uncertainty about a system or the information we lack to describe it, is a state quantity whose variation is independent of the path connecting two states. Thus, Brillouin’s law does not tell us where the energy cost of acquisition is paid in the operation. And this is inherent to the definition of what a state quantity is. Acquisition may not be a simple operation, but may consist of other, more elementary operations, which may differ depending on the specific data acquired. However, in the end, regardless of how the acquisition is performed, it will cost at least $T\ln 2$ on average per bit.
Brillouin’s law is a syllogism and states nothing more than its premises—in this case two mathematical results and the second law. It follows that its generality is as robust as that of the second law.
@@ III.3 Landauer’s Law on Data Erasure
Landauer’s competing argument [37] is based on two main ideas. The first is that “information is not a disembodied abstract entity; it is always tied to a physical representation. It is represented by engraving on a stone tablet, a spin, a charge, a hole in a punched card, a mark on paper, or some other equivalent” [39]. In other words, information is always supported by hardware with at least two discernible different configurations (two values: 0 or 1), namely a data bit. The second idea is that the acquisition of one bit of information necessarily passes by the erasure of its supporting data bit. The reasoning is as follows:
@@
    Any intelligent being has a finite memory; thus, the infinite cyclic acquisition of information about a dynamical system necessarily requires the erasure of data bits.
@@
    Erasing a data bit (a thermodynamical system) consists of setting it to an arbitrary value (say 0). The procedure must be able to work for a known or unknown initial value (i.e., it must be the same in both cases). This constraint automatically implies a two-step erasure process (see Figure 3):
@@
    Figure 3: Landauer erasure of a data bit consisting of a single particle in a two-compartment box. The procedure must be independent of the initial position of the particle: neither a simple shift of the particle towards 0 by means of two pistons keeping the phase-space volume constant, nor a quasi-static expansion with a single piston, is possible. According to Landauer, the first step is necessarily a thermodynamically irreversible free expansion.
@@
    The first step does not involve any exchanges with the environment, whereas the second dissipates at least $T\ln 2$ of heat, or equivalently, at constant internal energy (i.e., at constant temperature for a set of independent data bits), it requires a minimum work. The net balance of the two yields
@@
    which leads, as Brillouin’s law does (Equation (4)), to a zero energy balance for Szilard engine. The difference lies in the fact that, according to Landauer, erasure is a necessary step, and it is at this very point that the energy cost of data acquisition is paid.
@@
    Restricted context of data erasure: Does erasing one data bit really require at least $T\ln 2$ (Landauer’s limit) of work?
    To my knowledge, all the authors (see, e.g., [53, 30, 40, 26, 9, 18]) who have addressed this question actually used Landauer’s procedure that consists of (a) free expansion; (b) reversible compression. In this way, the authors simply test the second law rather than the novel aspect of Landauer’s idea; therefore, their results are unsurprising, as there is no error in Landauer’s calculation. What is new in Landauer’s assertion? It is the claim that there is no alternative erasing procedure. Thus, the only way to confirm Landauer’s law would be to search, in vain, for an alternative. However, this is not what was done. In reality, even within the restricted context of data erasure, several points of Landauer’s reasoning are questionable: the imperative to begin erasure with expansion, and the necessity of this expansion to be thermodynamically irreversible. Counterexamples have been provided [33, 34, 35], which invalidate the generality of Landauer’s procedure.
@@ IV Information Versus Data
The confusion between information and data also corresponds to the confusion between the loss (or deletion) of a bit of information and the erasure of a bit of data. The first refers to the observer’s knowledge, whereas the second only involves a physical object independent of them. A similar confusion in statistical mechanics has already led to inconsistencies. It is, therefore, not surprising to see others with Landauer’s idea. These inconsistencies make Landauer’s law incompatible with that of Brillouin—that is, incompatible with the triplet (thermodynamics–statistical mechanics–information theory).
@@ IV.1 Total Versus Incomplete Information
In thermodynamics, setting a data bit into the S state by removing the partition between two compartments containing a single particle (first step of Landauer erasure) is reversible or not, depending on whether the initial value of the bit is unknown or not [43] (see Figure 4). This problem is the same as that of the Gibbs paradox, and the solution is also the same: thermodynamics considers the case where the observer has incomplete information about the system, whereas Landauer, in his reasoning, implicitly assumes that the observer has all information at his disposal. There is no solution to this paradox other than those that reintroduce the observer, because in reality, taking these two cases into account (as Bennett does [3], a defender of Landauer’s thesis), already amounts to taking the observer into account.
Figure 4: Setting a data bit in state S by merging two compartments is thermodynamically irreversible (red path) if the initial value is known ($\Delta S=\ln 2$), but it is thermodynamically reversible (green path) if the initial value is unknown ($\Delta S=0$). This is reminiscent of the Gibbs paradox of mixing. Landauer’s viewpoint conflicts with thermodynamics.
In addition, the constraint of a single path for erasure, regardless of the initial value of the bit, stems from the requirement to work with unknown values. But this case is thermodynamically reversible. For known values, there is nothing preventing the use of two different reversible paths. The use of a single erasing procedure for known and unknown cases is an arbitrary choice by the observer. Here again lies a clear inconsistency between the intention of the theory and its result—the desire to eliminate the observer while reintroducing him.
@@ IV.2 Global Concept Versus Local Object
Information is a concept that must not be confused with its material embodiment. The proof is provided by Landauer himself: “Information […] is represented by engraving on a stone tablet, a spin, a charge, a hole in a punched card, a mark on paper, or some other equivalent” [39]. The material embodiment may change, but information remains the same. The denial of this difference leads to severe inconsistencies.
@@
A fundamental property of information is that it cannot be given twice. We can duplicate bits of data encoding certain information so that the storage space occupied is twice as large as before, but the corresponding amount of information remains unchanged. Conversely, if multiple copies of the same data bit are erased, the corresponding information will only be lost when the last copy is erased. To assert that information has been erased, one must take into account not only what happened locally, but also the existence or not of a copy somewhere in a more global space.
Information is a global concept, whereas a data bit is a local object.
@@
It could be argued that this is playing with words and that information can have different meanings. Actually, the common meaning of information as “The imparting of knowledge in general. Knowledge communicated concerning some particular fact, subject, or event” (Oxford Dictionary) is also that adopted by Shannon: if the language used for a message is known, the message can be further compressed, however indicating the language twice does not yield any additional space savings. This meaning is likewise shared by proponents of Landauer’s idea (“But what is information? A simple, intuitive answer is “what you don’t already know”” [40]. “If someone tells you that the earth is spherical, you surely would not learn much” [9]).
@@
The denial of the difference between information and data and the attempt to materialize information is culminating with the supposed “information-mass equivalence” [23, 25, 54, 55] that leads to paradoxes in relation to the corresponding mass deficit. For instance, in the framework of this equivalence, consider a data bit encoding a bit of information. Duplicate the data bit and make the original and the copy physically independent. Erase one of the two. The quantity of information remains unchanged, so the erased data bit does not display any mass deficit. Erase the second bit. Information is lost that corresponds to a mass deficit. How does the second bit “know” that the first has been erased? This contradicts the hypothesis of independence. But if we suppress this hypothesis and assert that independence is impossible, does that not mean that information is not a property of the local data bit, but a property of a global system?
@@
Yet another variation of this paradox. Consider a data bit encoding a bit of information whose initial value is 0. Erase the data bit (set its value to 0). There is no longer information stored by the data bit. Erase the bit again. What is the difference between the two erasures? If there is no difference, an eventual mass deficit should not be due to any loss of information. If there is a difference, how does the data bit “know” that it has already been erased? Only a global observer could.
@@
A data bit has a physical meaning at all scales—from that of a single particle in a two-compartment box, to the macroscopic scale. But information is meaningful only when it involves an observer; information is meaningful only at the macroscale. Information, even the smallest piece, is an emergent concept.